OpenStudy (anonymous):

Find (fog)(2) and (f+g)(2) when f(x)=1/x and g(x)=4x+9

OpenStudy (unklerhaukus):

\[(f\circ g)(x)=f\big(g(x)\big)\\ (f+g)(x)=f(x)+g(x)\]

OpenStudy (unklerhaukus):

can you find \(g(2)\) ?

OpenStudy (anonymous):

sorry I am still a bit confused

OpenStudy (unklerhaukus):

\[g(x)=4x+9\] \[g(2)=4(2)+9=4\times2+9=\]

OpenStudy (anonymous):

ok I kind of get it now, so I do the same for f(x)=1/x ?

OpenStudy (unklerhaukus):

Yeah, lets work out (f+g)(2) first because its a bit easier . What do you get?

OpenStudy (anonymous):